The length of a rectangle is 5 meters longer than the width. Find the length and width if the perimeter is 54 meters.
Answer by nadinadan(32)

Let's change the words problem in math operations.
The problem says:
L=5+W where L is the Length and W is the width.
And 2L+2W=54 The perimeter of the rectangle is L+L+W+W=2L+2W
So we have 2 equations:
L=5+W
2L+2W=54 we replace the first ecuation into the second one, so insted of L we will have (5+L)
So the second ecuation changes to
2(5+W)+2W=54 so 10+2W+2W=54 so 4W=54-10 so 4W=44 so W=11
than in L=5+W, we replace W=11, so L=5+11=16.
So L=16 and W=11.